An LC circuit includes a 0.045 µF capacitor and a 390 µH inductor. If the capacitor is rated at 600 V, what is the maximum inductor current that will keep the capacitor within its rating?
A
Answer #1

We know that, stored energy of a capacitor is

U = 1/2 C V^2

U = 1/2 x 0.045 x 10^-6 x 600^2 = 0.0081 J

that of an inductor = 1/2 L I^2

1/2 L I^2 = 0.0081

I = sqrt (2 x 0.0081 / L) = sqrt (2 x 0.0081/390 x 10^-6) = 6.45 A

Hence, I = 6.45 A